A nontoxic furniture polish can be made by combining vinegar and olive oil. The amount of oil should be three

times the amount of vinegar. How much of each ingredient is needed in order to make 38

oz of furniture​ polish?

Answer by rfer(16322)

18/4=4.5 vinegar
4.5*3=13.5 oil
--------
x+3x=18
4x=18
x=18/4
x=4.5
3x=13.5